A stone is thrown obliquely upwards with the throw angle 30degree relative to the horizontal plane. the highest point of the path after 0.15s. At what speed was the stone thrown
1
Expert's answer
2019-05-15T10:17:06-0400

The time of stone motion to the highest point

"t=\frac{v_0\sin\theta}{g}"

So, the initial speed

"v_0=\frac{gt}{\sin\theta}"

"=\frac{9.8\:\rm{m/s^2}\times 0.15\:\rm{s}}{\sin 30^{\circ}}=2.94\:\rm{m/s}"